# 1S-DAug: One-Shot Data Augmentation for Robust Few-Shot Generalization

### v1

@@ Abstract
Few-shot learning (FSL) challenges model generalization to novel classes based on just a few shots of labeled examples, a testbed where traditional test-time augmentations fail to be effective. We introduce 1S-DAug, a one-shot generative augmentation operator that synthesizes diverse yet faithful variants from just one example image at test time. 1S-DAug couples traditional geometric perturbations with controlled noise injection and a denoising diffusion process conditioned on the original image. The generated images are then encoded and aggregated, alongside the original image, into a combined representation for more robust FSL predictions. Integrated as a training-free model-agnostic plugin, 1S-DAug consistently improves FSL across standard benchmarks of 4 different datasets without any model parameter update, including achieving over 10% proportional accuracy improvement on the miniImagenet 5-way-1-shot benchmark. Codes will be released.
@@ 1 Introduction
Few-shot learning (FSL) is important for recognition systems deployed in the wild. While deep learning attains strong performance given abundant supervision, their accuracy degrades in rare-case generalization (Wang et al., 2020). Real-world data are long-tailed; rare categories with limited labels persist and cap overall system performance even as head classes continue to grow (Kang et al., 2020). Scarcity in the target domain induces a train-test gap that manifests as high generalization error on novel classes at test time (Wang et al., 2020). As both a testbed and a concrete instance of this long-tail problem, the FSL model must assign labels to previously unseen classes using only a handful of labeled examples per class (Vinyals et al., 2016; Snell et al., 2017). This low-data regime appears in practical settings such as medical imaging for rare diseases and autonomous driving with open-world, unpredictable events (Liu and Feng, 2024). The central challenge is to achieve robust generalization under label scarcity and distribution shift.
@@
  • We introduce 1S-DAug, a one-shot data augmentation method. Given a single image, it leverages attention-guided diffusion with controlled noise to generate faithful yet diverse variants suitable for inference on unseen classes. Though synthetic data augmentation with target-class supervision is not new, to the best of our knowledge, we are the first to focus on data augmentation given just one unlabelled image. Besides, 1S-DAug can fulfill the challenging FSL test-time augmentation requirement. As this direction is practically meaningful but underexplored, we analyse 1S-DAug’s integration with FSL.
  • We offer clean, intuitive theoretical analyses for 1S-DAug, where we mainly separate the FSL generalization risk to the empirical risk and the Rademacher complexity term. Under idealized assumptions, 1S-DAug improves both the empirical risk via diversity and the Rademacher complexity term under shrunk data radius. Though the risk bound is standard, its adaptation to FSL concisely captures the design rationale.
  • We implement 1S-DAug as a test-time plugin for trained FSL models, and we evaluate on benchmarks of four datasets, observing consistent gains (e.g., up to 20% relative improvement on 5-way-1-shot benchmarks) without any underlying parameter update to the off-the-shelf models, which is practically meaningful.
@@ 2 Related Work
For few-shot learning, augmentation-based approaches expand training diversity via feature hallucination (Hariharan and Girshick, 2017a) or GAN-driven synthesis (Wang et al., 2018), but they are largely training-time and rely on supervision from base classes. In contrast, test-time augmentation for FSL is harder; it must generate high-quality, class-faithful variants for unseen classes without any retraining or labels. The only explicit work in this direction is FSL-Rectifier (Bai et al., 2025), which uses FUNIT (Liu et al., 2019a) to combine the shape of one image with the class-defining style of another. This serves as a proof of concept, mainly targeting animal-face datasets with limited broad applicability. For diffusion models, attention-based conditioning adapters now inject external signals (e.g., an input image) into cross-attention, enabling faithful, controllable generation (Mou et al., 2023). SDEdit (Meng et al., 2022) adds controlled noise and then denoises with conditioning. However, lower noise level yields minimal change, and the higher sacrifices faithfulness (e.g., change object type), thus not suitable as data augmentation that requires both diversity and faithfulness. There are a few recent works using the diffusion-based synthetic images for training augmentation outside of few-shot learning, but they rely on fine-tuning based on text prompts or additional test-class samples, and are thus limited for test-time augmentation (He et al., 2023; Benigmim et al., 2023; Trabucco et al., 2024). More related works are available in Appendix Section §F.
@@ 5 Theoretical Analyses
We analyse 1S-DAug in the standard episodic few-shot setting with a single trainable encoder and a fixed Euclidean nearest-prototype classifier. We show that (i) a simple risk decomposition emphasises accuracy and diversity, and (ii) a tightened margin-based generalisation bound after augmentation, via both empirical margin and feature-radius reduction. Lastly, based on the bound, we compare training and test-time augmentation in Appendix B.6. Note that we focus on the overall theoretical logic instead of the details, which is simplifying but self-contained and sufficient to illustrate the design’s logical coherence.
5.1 Episodic Euclidean Model and Test-Time Augmentation
We consider the encoder-plus-Euclidean-prototype classifier of Section 4. In an $N$-way $K$-shot episode, class prototypes $p_{c}$ are formed by averaging support embeddings, and a query $q$ is assigned to the nearest prototype in squared Euclidean distance $\|\Phi_{\theta}(q)-p_{c}\|_{2}^{2}$. For the theory we reduce episodes to binary query-prototype pairs $x=(q,p)$ with label $y\in\{-1,1\}$, write
$\Omega_{\theta}(x):=\Phi_{\theta}(q)-p\quad\text{and}\quad g_{\theta}(x):=-\|\Omega_{\theta}(x)\|_{2}^{2},$
and assume a uniform radius bound $\|\Omega_{\theta}(x)\|_{2}\leq r_{0}$. Under test-time augmentation, multiple query views are combined by logit averaging; for squared Euclidean scores this is equivalent to using an averaged query embedding and hence an aggregated difference feature $\bar{\Omega}_{\theta}(x)$. Full episodic details and the logit-feature equivalence are deferred to Appendix B.1 and B.3.
@@ 5.2 Risk Decomposition into Accuracy and Diversity
We first quantify the ensemble effect at the pairwise level. For any real-valued predictor $g$ on pairs $x=(q,p)$ we use the scaled squared-loss risk
@@
Thus improvements come from (i) maintaining or improving single-view accuracy, and (ii) making the augmented predictions sufficiently diverse on hard examples. This matches the empirical behaviour of 1S-DAug, which is designed to generate plausible but different-shape query views.
@@ 5.3 Encoder Generalisation and Radius Reduction
In the pairwise reduction of §Section 5.1, each example is a query-prototype pair $x=(q,p)$ with label $y\in\{-1,+1\}$, difference feature $\Omega_{\theta}(x):=\Phi_{\theta}(q)-p$, and squared-Euclidean score $g_{\theta}(x):=-\|\Omega_{\theta}(x)\|_{2}^{2}$. We write the pairwise 0-11 risk as $R_{\mathrm{cls}}(\theta):=\mathbb{P}_{(x,y)\sim D}\bigl(y\,g_{\theta}(x)\leq 0\bigr),$ and assume the uniform pre-augmentation radius bound $\|\Omega_{\theta}(x)\|_{2}\leq r_{0}$. Under test-time augmentation, query logit averaging is equivalent (up to a class-independent constant) to nearest-prototype classification using the averaged query embedding (Section 5.1, formalized in Appendix B.3); hence we may analyze the aggregated feature $\bar{\Omega}_{\theta}(x)$ and aggregated score $\bar{g}_{\theta}(x):=-\|\bar{\Omega}_{\theta}(x)\|_{2}^{2}$.
@@
Let $u_{\theta}(x):=y\,g_{\theta}(x)$ denote the signed margin and let $\tau_{\rho}$ be the standard $\rho$-margin surrogate (Appendix B.2). For a sample $S=\{(x_{i},y_{i})\}_{i=1}^{m}$, define the empirical margin risk $\widehat{R}_{S,\rho}(\theta):=\frac{1}{m}\sum_{i=1}^{m}\tau_{\rho}(u_{\theta}(x_{i}))$ and the empirical Rademacher complexity $\widehat{\mathfrak{R}}_{S}(\mathcal{G})$ of the score class $\mathcal{G}:=\{g_{\theta}:\theta\in\Theta\}$. The following is a standard contraction-based margin bound (proof in Appendix B.2).
@@
For any $\rho>0$ and $\delta\in(0,1)$, with probability at least $1-\delta$ over $S\sim D^{m}$, every $\theta$ satisfies
$R_{\mathrm{cls}}(\theta)\;\leq\;\widehat{R}_{S,\rho}(\theta)+\frac{2}{\rho}\widehat{\mathfrak{R}}_{S}(\mathcal{G})+\sqrt{\frac{\log(1/\delta)}{2m}}.$ (13)
To make the complexity term explicit for Euclidean scores, we control $\widehat{\mathfrak{R}}_{S}(\mathcal{G})$ via the encoder Lipschitz constant and the feature radius (proof sketch in Appendix B.2).
@@
Assume $\Phi_{\theta}$ is realised by a feedforward network with $1$-Lipschitz nonlinearities and layer spectral norms whose product is bounded by $L_{\mathrm{enc}}$, and assume a uniform radius bound $\|\Omega_{\theta}(x)\|_{2}\leq r$ for all $x$ and $\theta$. Then there exists an architecture-dependent constant $C_{\mathrm{enc}}>0$ such that $\widehat{\mathfrak{R}}_{S}(\mathcal{G})\;\leq\;C_{\mathrm{enc}}\,L_{\mathrm{enc}}\,\frac{r}{\sqrt{m}}.$
Combining Theorem 5.2 and Lemma 5.3 yields a bound in which the complexity term scales linearly with the effective feature radius. This is the mechanism by which aggregation can tighten generalisation.
Why Test-time Aggregation Works.
Let $\widehat{S}$ denote the same labeled pairs as $S$ but evaluated with aggregated difference features $\bar{\Omega}_{\theta}(x_{i})$ and aggregated score $\bar{g}_{\theta}$. Applying Theorem 5.2 to $\bar{g}_{\theta}$ yields the same bound with $\widehat{R}_{S,\rho}(\theta)$ replaced by $\widehat{R}_{\widehat{S},\rho}(\theta)$ and the radius parameter $r$ replaced by an aggregated radius $\hat{r}$ such that $\|\bar{\Omega}_{\theta}(x)\|_{2}\leq\hat{r}$. Since Lemma 5.3 is linear in $r$, any contraction $\hat{r}<r_{0}$ directly contracts the complexity term. A simple stability statement relates the empirical margin risks before and after aggregation (proof in Appendix B.4).
Proposition 5.4 (Probabilistic radius reduction with $M$ augmentations).
Let $\{\Omega_{i}^{(k)}\}_{i=1,k=0}^{m,\;M}$ be i.i.d. from a continuous distribution and $\widehat{\Omega}_{i}:=\frac{1}{M+1}\sum_{k=0}^{M}\Omega_{i}^{(k)}$. Writing $R_{\mathrm{orig}}^{\max}:=\max_{i}\|\Omega_{i}^{(0)}\|_{2}$ and $\widehat{R}^{\max}:=\max_{i}\|\widehat{\Omega}_{i}\|_{2}$, we have $\mathbb{P}\bigl(\widehat{R}^{\max}<R_{\mathrm{orig}}^{\max}\bigr)\;\geq\;\frac{1}{M+1}.$
Proof is in Appendix B.5. Taken together, Lemma 5.3 and Proposition 5.4 illustrate that the multi-view aggregation can tighten the complexity term through radius contraction. Additionally, the potentially reduced risk empowered by both accuracy and diversity as specified in Section §5.2 also indicates lower empirical risk compared to before the augmentation. Therefore, the overall generalization bound in 5.2 tends to tighten.
@@
Training-time augmentation on base classes primarily tightens a bound on the base risk by increasing sample size, but any transfer to novel classes typically incurs an additional distribution-shift discrepancy term. In contrast, 1S-DAug targets the novel classes directly. More discussions are available in Appendix B.6.
@@ 6 Experiments
Table 1/2/3 summarizes 5-way-1-shot and 5-way-5-shot results with Res12/ConvNet/ViTSmall/SwinTiny backbones, and our method with 1/2/3 additional augmentations are denoted as 1S-DAug-1/2/3 respectively. Note that we directly adopt 5-way-1-shot FSL models for the 5-way-5-shot evaluation, and Table 2 contains 5-way-1-shot results on CUB and Animals. As reported, test-time 1S-DAug consistently improves over the corresponding non-augmented baselines and over prior strong Res12 methods reported under the same backbone (e.g., on miniImagenet, FEAT improves from $63.31\%$ to $69.25\%$, a maximum absolute gain of +5.94+5.94 percentage points, which achieves the highest among all the reported FSL works with Res12 backbones). Note that our method performs significantly better on the Animals dataset compared to the prior work that uses GAN reconstruction, which only achieves 57.90% on the 5-way-1-shot setup with one augmentation (Bai et al., 2025). The limitation GAN generation quality has been discussed in (Bai et al., 2025; Liu et al., 2019a), and the related failure case visualization is available in Appendix Figure 4. All our gains persist across 4 datasets (e.g., miniImagenet improves by at least $2.89\%$, tieredImagenet by $3.88\%$, CUB by $9.12\%$, and Animals by $1.23\%$ on the 5-way-1-shot benchmarks), implying a high probability for: (i) the image-conditioned diffusion step preserves class-defining content sufficiently, and (ii) the shape tweak with noising creates diversity without compromising faithfulness. Additionally, ViTSmall/SwinTiny backbones are pretrained on ImageNet-1k (Russakovsky et al., 2015), and we still achieve consistent non-trivial improvement over these more knowledgeable backbones.
@@ 7 Limitations and Future Work
Our approach has two main implementation limitations. First, the augmentation introduces inference overhead (See Appendix C). Reducing computational cost (e.g., via lighter generative backbones or faster denoising schedules) is useful, but the overhead can automatically attenuate as computation infrastructure develops. Second, we use the Stable-Diffusion-v1.5 generator (Rombach et al., 2022b)), where we partially go beyond the FSL training data split. However, the FSL training splits are generally not enough to train a good-performing stable diffusion model, while the standard pretrained models are already available. Therefore, choosing the pretrained model is natural in both the real-world and research set-up. Also, we freeze all pretrained models and do not fine-tune on the test set. Importantly, our main contribution is not the specific implementation, but the algorithm design that scales with computation development.

### v3

@@ Abstract
Few-shot learning (FSL) challenges model generalization to novel classes based on just a few shots of labeled examples, a testbed where traditional test-time augmentations fail to be effective. We introduce 1S-DAug, a one-shot generative augmentation operator that synthesizes diverse yet faithful variants from just one example image at test time. 1S-DAug couples traditional geometric perturbations with controlled noise injection and a denoising diffusion process conditioned on the original image. The generated images are then encoded and aggregated, alongside the original image, into a combined representation for more robust FSL predictions. Integrated as a training-free model-agnostic plugin, 1S-DAug consistently improves FSL across standard benchmarks of 4 different datasets without any model parameter update, including achieving up to 20% relative accuracy improvement on the miniImagenet 5-way-1-shot benchmark. Code will be released.
@@ 1 Introduction
Few-shot learning (FSL) is important for recognition systems deployed in the wild. While deep learning attains strong performance given abundant supervision, their accuracy degrades in rare-case generalization (Wang et al., 2020). Real-world data are long-tailed; rare categories with limited labels persist and cap overall system performance even as head classes continue to grow (Kang et al., 2020). Scarcity in the target domain induces a train-test gap that manifests as high generalization error on novel classes at test time (Wang et al., 2020). As both a testbed and a concrete instance of this long-tail problem, the FSL model needs to assign labels to previously unseen classes using only a handful of labeled examples per class (Vinyals et al., 2016; Snell et al., 2017). This low-data regime appears in practical settings such as medical imaging for rare diseases and autonomous driving with open-world, unpredictable events (Liu and Feng, 2024). The central challenge is to achieve robust generalization under label scarcity and distribution shift.
@@
  • We introduce 1S-DAug, a one-shot data augmentation method. Given a single image, it leverages attention-guided diffusion with controlled noise to generate faithful yet diverse variants suitable for inference on unseen classes. Though synthetic data augmentation with target-class supervision is not new, data augmentation given just one unlabelled image is underexplored. Besides, 1S-DAug can fulfill the challenging FSL test-time augmentation requirement. As this direction is practically meaningful but also underexplored, we analyse 1S-DAug’s integration with FSL.
  • We implement 1S-DAug as a test-time plugin for trained FSL models, and we evaluate on benchmarks of four datasets, observing consistent gains (e.g., up to 20% relative improvement on 5-way-1-shot benchmarks) without any underlying parameter update to the off-the-shelf models.
  • We offer intuitive analyses for why 1S-DAug practically works, where we mainly separate the FSL generalization risk to the empirical risk and the Rademacher complexity term. Under idealized assumptions, 1S-DAug improves both the empirical risk via ensemble, and the reduced Rademacher complexity term under shrunk data radius. The risk bound adaptation to FSL concisely captures the design rationale.
@@ 2 Related Work
For few-shot learning, augmentation-based approaches expand training diversity via feature hallucination (Hariharan and Girshick, 2017a) or GAN-driven synthesis (Wang et al., 2018), but they are largely training-time and rely on supervision from base classes. In contrast, test-time augmentation for FSL is harder; it needs to generate high-quality, class-faithful variants for unseen classes without any retraining or labels. The only explicit work in this direction is FSL-Rectifier (Bai et al., 2025), which uses FUNIT (Liu et al., 2019a) to combine the shape of one image with the class-defining style of another. This serves as a proof of concept, mainly targeting animal-face datasets with limited broad applicability. For diffusion models, attention-based conditioning adapters now inject external signals (e.g., an input image) into cross-attention, enabling faithful, controllable generation (Mou et al., 2023). SDEdit (Meng et al., 2022) adds controlled noise and then denoises with conditioning. However, lower noise level yields minimal change, and the higher sacrifices faithfulness (e.g., change object type), thus not suitable as data augmentation that requires both diversity and faithfulness. There are a few recent works using the diffusion-based synthetic images for training augmentation outside of few-shot learning, but they rely on fine-tuning based on text prompts or additional test-class samples, and are thus limited for test-time augmentation (He et al., 2023; Benigmim et al., 2023; Trabucco et al., 2024). More related works are available in Appendix Section §F.
@@ 4 Method
Equivalence of the logit-averaging and feature-averaging is discussed in Appendix B.2.
@@ 5 Theoretical Analyses
We analyse 1S-DAug in the standard episodic few-shot setting with a single trainable encoder and a fixed Euclidean nearest-prototype classifier. We show that (i) a simple risk decomposition emphasises accuracy and diversity, and (ii) a tightened margin-based generalisation bound after augmentation, via both empirical margin and feature-radius reduction. Lastly, based on the bound, we heuristically compare training and test-time augmentation in Appendix B.5.
@@ 5.1 Risk Decomposition into Accuracy and Diversity
We first illustrate the ensemble effect intuition at the pairwise level. For any real-valued predictor $g$ on pairs $x=(q,p)$ we use the scaled squared-loss risk
@@
Thus improvements come from (i) maintaining or improving single-view accuracy, and (ii) augmentation diversity. This matches the 1S-DAug design for plausible but different-shape sample views.
5.2 Episodic Euclidean Model and Test-Time Augmentation
We consider the encoder-plus-Euclidean-prototype classifier. In an $N$-way $K$-shot episode, class prototypes $p_{c}$ are formed by averaging support embeddings, and a query $q$ is assigned to the nearest prototype in squared Euclidean distance $\|\Phi_{\theta}(q)-p_{c}\|_{2}^{2}$. For the theory we reduce episodes to binary query-prototype pairs $x=(q,p)$ with label $y\in\{-1,1\}$, write
$\Omega_{\theta}(x):=\Phi_{\theta}(q)-p\quad\text{and}\quad g_{\theta}(x):=\beta-\|\Omega_{\theta}(x)\|_{2}^{2},$
for a fixed threshold $\beta\in\mathbb{R}$. This mechanism differs slightly from Section 4, acting as a simplified proxy of analysis.
@@ 5.3 Encoder Generalisation and Radius Reduction
Let $u_{\theta}(x):=y\,g_{\theta}(x)$ denote the signed margin and let $\tau_{\rho}$ be the standard $\rho$-margin surrogate (Appendix B.3). For a generic pairwise sample $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{m}$, define the empirical margin risk $\widehat{R}_{\mathcal{T},\rho}(\theta):=\frac{1}{m}\sum_{i=1}^{m}\tau_{\rho}(u_{\theta}(x_{i}))$ and the empirical Rademacher complexity $\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})$ of the score class $\mathcal{G}:=\{g_{\theta}:\theta\in\Theta\}$. We have a contraction-based margin bound (proof in Appendix B.3):
@@
For any $\rho>0$ and $\delta\in(0,1)$, with probability at least $1-\delta$ over $\mathcal{T}\sim D^{m}$, every $\theta$ satisfies
$R_{\mathrm{cls}}(\theta)\;\leq\;\widehat{R}_{\mathcal{T},\rho}(\theta)+\frac{2}{\rho}\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})+\sqrt{\frac{\log(1/\delta)}{2m}}.$ (13)
To make the complexity term explicit for Euclidean scores, we control $\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})$ via the encoder Lipschitz constant and the feature radius (proof in Appendix B.3):
@@
Assume $\Phi_{\theta}$ is realised by a feedforward network with $1$-Lipschitz nonlinearities and layer spectral norms whose product is bounded by $L_{\mathrm{enc}}$, assume $B_{X}:=\sup_{q\in\mathcal{X}}\|q\|_{2}<\infty$, assume for simplicity that the prototypes are fixed, and assume a uniform radius bound $\|\Omega_{\theta}(x)\|_{2}\leq r$ for all $x$ and $\theta$. Then there exists an architecture-dependent constant $C_{\mathrm{enc}}>0$ such that $\widehat{\mathfrak{R}}_{\mathcal{T}}(\mathcal{G})\;\leq\;C_{\mathrm{enc}}\,B_{X}\,L_{\mathrm{enc}}\,\frac{r}{\sqrt{m}}.$
Applying Theorem 5.2 to test-time augmentation yields the same bound with both the empirical risk and Rademacher complexity terms shrunk with high probability. In particular, the empirical risk tends to shrink due to the ensemble effect (Zhou, 2012). The $B_{X}$ and $r$ factors in the Rademacher complexity term tend to shrink with high-probability feature radius reduction specified in Proposition 5.4 below:
Proposition 5.4 (Radius reduction with one augmentation).
Let $\{\Omega_{i}^{(0)},\Omega_{i}^{(1)}\}_{i=1}^{m}$ be i.i.d. from a continuous distribution and define $\bar{\Omega}_{i}:=\frac{1}{2}\bigl(\Omega_{i}^{(0)}+\Omega_{i}^{(1)}\bigr).$ Writing $r_{\mathrm{orig}}^{\max}:=\max_{i}\|\Omega_{i}^{(0)}\|_{2}$ and $\hat{r}^{\max}:=\max_{i}\|\bar{\Omega}_{i}\|_{2}$ and noting $m>1$ for few-shot classification, we have $\mathbb{P}\bigl(\hat{r}^{\max}<r_{\mathrm{orig}}^{\max}\bigr)>0.5.$
Averaging more than one augmented view is expected to further stabilize the aggregated feature radius, but we use only the one-augmentation statement above.
@@
Training-time augmentation on base classes primarily tightens a bound on the base risk by increasing sample size, but any transfer to novel classes typically incurs an additional distribution-shift discrepancy term. In contrast, 1S-DAug targets the novel classes directly. Heuristic discussions are available in Appendix B.5.
@@ 6 Experiments
Table 1/2/3 summarizes 5-way-1-shot and 5-way-5-shot results with Res12/ConvNet/ViTSmall/SwinTiny backbones, and our method with 1/2/3 additional augmentations are denoted as 1S-DAug-1/2/3 respectively. Note that we adopt 5-way-1-shot FSL models for the 5-way-5-shot evaluation, and Table 2 contains 5-way-1-shot results on CUB and Animals. As reported, test-time 1S-DAug consistently improves over the corresponding non-augmented baselines and over prior strong Res12 methods reported under the same backbone (e.g., on miniImagenet, FEAT improves from $63.31\%$ to $69.25\%$, a maximum absolute gain of +5.94+5.94 percentage points, which achieves the highest among all the reported FSL works with Res12 backbones). Note that our method performs significantly better on the Animals dataset compared to the prior work that uses GAN reconstruction, which only achieves 57.90% on the 5-way-1-shot setup with one augmentation (Bai et al., 2025). The limitation GAN generation quality has been discussed in (Bai et al., 2025; Liu et al., 2019a), and the related failure case visualization is available in Appendix Figure 4. All our gains persist across 4 datasets (e.g., miniImagenet improves by at least $2.89\%$, tieredImagenet by $3.88\%$, CUB by $9.12\%$, and Animals by $1.23\%$ on the 5-way-1-shot benchmarks), implying that: (i) the image-conditioned diffusion step preserves class-defining content sufficiently, and (ii) the shape tweak with noising creates diversity without compromising faithfulness. Additionally, ViTSmall/SwinTiny backbones are pretrained on ImageNet-1k (Russakovsky et al., 2015), and we still achieve consistent non-trivial improvement over these more knowledgeable backbones.
@@ 7 Limitations and Future Work
Our approach has two main implementation limitations. First, the augmentation introduces inference overhead (See Appendix C). Reducing computational cost (e.g., via lighter generative backbones or faster denoising schedules) is useful, but the overhead can automatically attenuate as computation infrastructure develops. Second, we use the Stable-Diffusion-v1.5 generator (Rombach et al., 2022b)), where we partially go beyond the FSL training data split. The FSL training splits are generally not enough to train a good-performing stable diffusion model. However, it is still meaningful to explore how the standard pretrained models can help with FSL, as they become more and more mature and popular. Therefore, choosing the pretrained model is natural in both the real-world and research set-up. Also, we freeze all pretrained models and do not fine-tune on the test set. Importantly, our main contribution is not the specific implementation, but the algorithm design that scales with computation development.